### v2

# Márcia Lemos-Silva1

Abstract
Based on the classical continuous system initially proposed by Bailey in 1975, we present a novel Susceptible–Infected–Removed (SIR) model defined in quantum time, where the temporal evolution is governed by a non-uniform time grid. An explicit analytical solution is derived, and the long-term behavior of the susceptible, infected, and removed individuals is analyzed. Moreover, we prove the model preserves dynamic consistency with its continuous counterpart, as evidenced by the non-negativity of solutions and their corresponding qualitative agreement with the continuous dynamics. All results are further supported by illustrative examples.
Keywords: quantum SIR model, exact solution, asymptotic behavior.
MSC 2020: 81Q80 (primary), 92D30 (secondary).
1 Introduction
The Susceptible–Infected–Removed (SIR) model is well-established and has been extensively studied over the years [3]. Despite its long history, it continues to be of current great interest, since it provides a solid foundation for understanding and predicting the spread of diseases within populations [6, 16, 17].
In [1], Norman Bailey proposed the following SIR model:
$\begin{cases}x^{\prime}=-\dfrac{bxy}{x+y},\\ y^{\prime}=\dfrac{bxy}{x+y}-cy,\\ z^{\prime}=cy,\end{cases}$ (1)
where $b,c\in\mathbb{R}^{+}$, and where $x(t_{0})=x_{0}$, $y(t_{0})=y_{0}$, and $z(t_{0})=z_{0}$ are non-negative and, moreover, $x,y,z:\mathbb{R}\rightarrow\mathbb{R}^{+}_{0}$. This model adopts a well-known compartmental approach, wherein three variables represent different groups within the population. The variable $x$ represents the proportion of susceptible individuals, $y$ represents the proportion of infected individuals, and $z$ represents the proportion of removed individuals. By adding the right-hand side of equations (1), one has $x^{\prime}(t)+y^{\prime}(t)+z^{\prime}(t)=0$, which is equivalent to $x(t)+y(t)+z(t)=N$, for all $t$, where the constant $N:=x_{0}+y_{0}+z_{0}$ denotes the total population under study. The exact solution of (1) was found in [5]. More recently, a different method was proposed to solve (1), generalizing the result of [5] to non-autonomous cases [2].
Quantum calculus, also known as $q$-calculus, is a branch of mathematics that extends the traditional calculus to a framework that incorporates $q$-analogues, that is, mathematical structures that replace certain variables with $q$-deformations, where $q$ is a parameter that controls the deformation [8, 9]. In the context of quantum systems or quantum phenomena, mathematical modeling often requires specialized mathematical frameworks that can account for unique properties of quantum mechanics. Mathematical models built within quantum calculus can be used to study a wide range of phenomena, such as quantum entanglement, quantum information, quantum statistical mechanics, and quantum field theory. These models aid in understanding the behavior of quantum systems, predicting their properties.
Here, we propose and investigate a SIR model formulated in quantum time. One of the main motivations for considering such a model lies in the non-uniform structure of the quantum time scale. Unlike the standard discrete-time framework, where time progresses in uniform steps, quantum time evolves on a increasing step. This means that during the early stages of an epidemic, where dynamics tend to be more sensitive, the step is finer, turning to a coarser one as the system stabilizes. Quantum modeling offers a more natural discretization for many real-world phenomena, where the initial dynamics changes rapidly before approaching an equilibrium. To the best of our knowledge, no SIR model has yet been formulated in quantum time, making our approach a novel contribution to the field. Thus, regarding applications, the quantization of infectious disease modeling may effectively address some limitations of traditional continuous compartmental models by introducing a novel approach to understand disease dynamics. Furthermore, unlike many SIR models, our proposed model has an exact solution. Finally, we prove that all its solutions remain non-negative and that their long-term behavior are coherent with (1), which guarantees consistency.
To describe the dynamics of the model, we make use of the $q$-derivatives defined in quantum calculus as
$D_{q}f(x)=\frac{f(qx)-f(x)}{(q-1)x},$ (2)
where $D_{q}f$ denotes the $q$-derivative of function $f$. Moreover, $f(x)$ and $f(qx)$ represent the value of the function at $x$ and $qx$, respectively, and (q1)x(q-1)x denotes the step size that, as usual, measures the interval between $x$ and $qx$. Note that, when $q$ approaches 1, the $q$-derivative recovers its classical counterpart.
The paper is organized as follows. In Section 2, we introduce the SIR model using $q$-derivatives and derive the explicit $q$-model. Since our goal is to prove dynamic consistency, we begin by proving the non-negativity and boundedness of solutions. In the same section, we find the equilibrium points and derive the exact solution. Section 2 ends with the asymptotic behavior of the solution. We end our work with Section 3, providing some examples that illustrate the obtained results.
2 Main Results
To obtain the quantum system, we extend the Mickens method [10] to the framework of quantum calculus [8, 9]. Such method is a Nonstandard Finite Difference (NSFD) scheme that, unlike classical methods, such as Euler and Runge–Kutta, ensures dynamic consistency [13]. Developed by Mickens, this method presents some rules specifically designed to avoid the instabilities and inconsistencies common in standard discretization methods [10, 11, 12]. This method has already been successfully applied to various systems, e.g., [4, 7, 14, 15]. However, in our work, since we are using the concept of $q$-derivatives for the approximation of derivatives, we follow only one of Mickens’ rules, which states that the linear and nonlinear terms of the system may need to be replaced by nonlocal terms. As we shall see, this rule will be crucial for proving the non-negativity and boundedness of solutions.
Consider the quantum SIR model of the form
$\begin{cases}D_{q}x(t)=-\dfrac{bx(qt)y(t)}{x(t)+y(t)},\\ D_{q}y(t)=\dfrac{bx(qt)y(t)}{x(t)+y(t)}-cy(qt),\\ D_{q}z(t)=cy(qt),\end{cases}$ (3)
where $D_{q}x(t)$, $D_{q}y(t)$ and $D_{q}z(t)$ represent the $q$-derivatives with respect to $x$, $y$ and $z$, respectively, and $qt$ denotes the moment following $t$. Moreover, $x(t_{0})>0$, $y(t_{0})>0$ and $z(t_{0})\geq 0$, and $x,y:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}_{0}$, $z:\mathbb{R}^{+}_{0}\rightarrow\mathbb{R}^{+}_{0}$, $b,c\in\mathbb{R}^{+}$ and $q>1$. To compare our model (3) with the ones available in the literature one can take the limit when $q$ tends to one from the right. In the limit, when $q\rightarrow 1^{+}$, model (3) converges to the classical continuous SIR model (1).
Substituting (2) in (3), we obtain (3) in the form
$\begin{cases}x(qt)=\dfrac{x(t)(x(t)+y(t))}{x(t)+y(t)(1+b(q-1)t)},\\ y(qt)=\dfrac{y(t)(1+b(q-1)t)(x(t)+y(t))}{(1+c(q-1)t)(x(t)+y(t)(1+b(q-1)t))},\\ z(qt)=\dfrac{c(q-1)ty(t)(1+b(q-1)t)(x(t)+y(t))}{(1+c(q-1)t)(x(t)+y(t)(1+b(q-1)t))}+z(t).\end{cases}$ (4)
2.1 Non-negativity and boundedness of solutions
Ensuring the non-negativity of solutions and maintaining their boundedness are both crucial properties in epidemiological models. Follows our first result.
Proposition 1.
If the initial conditions are non-negative and the parameter values of the system (4) are positive, then all the solutions remain positive for all $t>t_{0}$. Moreover, the quantum model (3) guarantees that the population remains constant over time.
Proof.
Let $(x(t_{0}),y(t_{0}),z(t_{0}))\in\mathbb{R}^{3}_{+}$. Since all parameters are positive and, by definition, $q>1$, then all equations of system (4) are positive. Thus, the non-negativity condition is trivially satisfied. Now, let the total population be defined as $N(t)=x(t)+y(t)+z(t)$. Adding the right-hand side of equations (3), we have $D_{q}N(t)=0$, which means, by the properties of the quantum derivative [8, 9], that the population $N(t)$ remains constant along time $t$. ∎
2.2 Equilibrium points
To have dynamical consistency, the equilibrium points of the quantum model have to be the same of the continuous model (1).
Equating all the $q$-derivatives to zero, it follows easily that the equilibrium points of system (3) are given by $(\alpha,0,N-\alpha)$, with $\alpha\geq 0$. This is coherent with the equilibrium points of the continuous-time model (1) stated in [1].
2.3 Exact solution
Here, we derive the exact solution of system (4). Note that, since $N=x(qt)+y(qt)+z(qt)$ for all $t$, it means it is enough to solve the first two equations of system (4): after we know $x(t)$ and $y(t)$, we can immediately compute $z(qt)$ using the equality z(qt)=Nx(qt)y(qt)z(qt)=N-x(qt)-y(qt).
Remark 1.
In the following results, we use the notation $x(t_{n})$, where $n\in\mathbb{N}_{0}$. By definition, $t_{1}=qt_{0}$, $t_{2}=qt_{1}=q^{2}t_{0}$, …, i.e., $t_{n}=q^{n}t_{0}$.
Remark 2.
Let $x_{0},x_{1},x_{2},\dots$ be a sequence. Then, by convention, the product $\prod_{i=n}^{m}x_{i}$ is equal to 1 when $n>m$.
Theorem 1.
The exact solution of system (4) is given by
$\begin{cases}x(t_{n})=\displaystyle\frac{x(t_{0})\left(\bar{\kappa}+1\right)}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}},\\ \\ y(t_{n})=\frac{y(t_{0})}{\xi_{0}}\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\left(\frac{1}{\xi_{i+1}}\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}\right),\\ \\ z(t_{n})=N-\left(x(t_{0})+\frac{y(t_{0})}{\xi_{0}}\displaystyle\prod_{i=0}^{n-2}\left(\frac{1}{\xi_{i+1}}\right)\right)\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}},\end{cases}$ (5)
where $\bar{\kappa}=\frac{x(t_{0})}{y(t_{0})}$ and $\xi_{k}=\frac{1+c(q-1)q^{k}t_{0}}{1+b(q-1)q^{k}t_{0}}$, $k\in\{0,1,\dots,n-1\}$.
Proof.
We do the proof by induction. For $n=1$, we have
$x(t_{1})$ $=x(t_{0})\cdot\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}$ (6)
$=\frac{x(t_{0})(x(t_{0})+y(t_{0}))}{x(t_{0})+y(t_{0})(1+b(q-1)t_{0})},\quad\text{by the definition of}\,\bar{\kappa},$
and
$y(t_{1})$ $=\frac{y(t_{0})}{\xi_{0}}\cdot\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}$ (7)
$=\frac{y(t_{0})(1+b(q-1)t_{0})(x(t_{0})+y(t_{0}))}{(1+c(q-1)t_{0})(x(t_{0})+y(t_{0})(1+b(q-1)t_{0}))},\quad\text{by the definition of}\,\bar{\kappa}\,\text{and}\,~\xi_{0}.$
Both (6) and (7) are true from system (4). Now, let us state the inductive hypothesis that (5) holds true for a certain $n=m$. We want to prove that it remains valid for $n=m+1$. Thus,
$x(t_{m+1})$ $=x(t_{0})\cdot\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\prod_{i=0}^{m-1}\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}$
$=x(t_{0})\cdot\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\prod_{i=0}^{m-2}\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}$
$\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}}{1+b(q-1)q^{m}t_{0}+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}}$
$=x(t_{m})\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}}{1+b(q-1)q^{m}t_{0}+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}},\,\text{by inductive hypothesis}.$
Now, note that the first two equations of (5), that hold true by the inductive hypothesis, can be rewritten as
$x(t_{0})=\frac{x(t_{n})}{\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}}$ (8)
and
$y(t_{0})=\frac{y(t_{n})\xi_{0}}{\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\left(\frac{1}{\xi_{i+1}}\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}\right)}.$ (9)
Using (8) and (9), and the definition of $\bar{\kappa}$, one can write that
$\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}$ $=\frac{x(t_{m})\prod_{i=0}^{m-2}\frac{1}{\xi_{i+1}}}{y(t_{m})\xi_{0}}\cdot\prod_{j=0}^{m-1}\xi_{j}$ (10)
$=\frac{x(t_{m})}{y(t_{m})\prod_{i=0}^{m-1}\xi_{i}}\cdot\prod_{j=0}^{m-1}\xi_{j}$
$=\frac{x(t_{m})}{y(t_{m})}.$
Thus, applying (10), $x(t_{m+1})$ becomes
$x(t_{m+1})$ $=x(t_{m})\cdot\frac{x(t_{m})+y(t_{m})}{x(t_{m})+y(t_{m})(1+b(q-1)q^{m}t_{0})}$ (11)
$=x(t_{m})\cdot\frac{x(t_{m})+y(t_{m})}{x(t_{m})+y(t_{m})(1+b(q-1)t_{m})},\quad\text{because}\,q^{m}t_{0}=t_{m}.$
Moreover,
$y(t_{m+1})$ $=y(t_{0})\cdot\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\prod_{i=0}^{m-1}\left(\frac{1}{\xi_{i+1}}\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}\right)$
$=y(t_{0})\cdot\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\prod_{i=0}^{m-2}\left(\frac{1}{\xi_{i+1}}\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}\right)$
$\cdot\frac{1}{\xi_{m}}\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}}{1+b(q-1)q^{m}t_{0}+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}}$
$=\frac{y(t_{m})}{\xi_{m}}\cdot\frac{1+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}}{1+b(q-1)q^{m}t_{0}+\bar{\kappa}\prod_{j=0}^{m-1}\xi_{j}},\,\text{by inductive hypothesis}.$
Finally, given that $q^{m}t_{0}=t_{m}$, and applying (10) and the definition of $\xi_{m}$, it follows that
$y(t_{m+1})=\frac{y(t_{m})(1+b(q-1)t_{m})(x(t_{m})+y(t_{m}))}{(1+c(q-1)t_{m})(x(t_{m})+y(t_{m})(1+b(q-1)t_{m}))}.$ (12)
We end by noting that both (11) and (12) are also in agreement with system (4). Furthermore, since $N=x(t_{n})+y(t_{n})+z(t_{n})$, we have
$z(t_{n})=N-\left(x(t_{0})+\frac{y(t_{0})}{\xi_{0}}\displaystyle\prod_{i=0}^{n-2}\left(\frac{1}{\xi_{i+1}}\right)\right)\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\frac{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}{1+b(q-1)q^{i+1}t_{0}+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}.$
The proof is complete. ∎
Note that $x(t_{n})$ and $y(t_{n})$ of (5) can be rewritten as
$x(t_{n})=x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}},$ (13)
and
$y(t_{n})=\frac{y(t_{0})}{\xi_{0}}\cdot\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\left(\frac{1}{\xi_{i+1}}\cdot\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}}\right).$ (14)
The reproduction number $\mathcal{R}_{0}$ is one of the most important concepts in epidemiology. For system (1), it is well-known that
$\mathcal{R}_{0}=\frac{b}{c}.$
Since the underlying principles of disease transmission and recovery do not change, the expression for $\mathcal{R}_{0}$ remains valid for the quantum time system (3) or (4). In the following results, the focus is on the asymptotic stability of the equilibrium points considering the value of $\mathcal{R}_{0}$.
Theorem 2.
Assume that
1ξi+111+b(q1)qi+1t01+κ¯j=0iξj<1,for all
i
0
[…]
If $\mathcal{R}_{0}\geq 1$, then all solutions (5) of system (4) converge to the equilibrium point $(0,0,N)$.
Proof.
First, let us consider $\mathcal{R}_{0}=1$, that is, $b=c$. Then,
$\xi_{j}=\frac{1+c(q-1)q^{j}t_{0}}{1+b(q-1)q^{j}t_{0}}=1.$
This means that
$x(t_{n})=x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}}}.$
Clearly, $\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}<1$ and $\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}}}<1$ and, this way, one can easily conclude that $\underset{n\to\infty}{\lim}x_{n}=0$. The same inference can be drawn regarding $y(t_{n})$. Moreover, since z(t)=Nx(t)y(t)z(t)=N-x(t)-y(t), it is clear that under these circumstances all solutions of (5) converge to $(0,0,N)$.
Let us now consider the remaining case where $\mathcal{R}_{0}>1$, that is, $b>c$. This leads to $\xi_{j}<1$. Considering (13), we can state $a_{i}$ as
$a_{i}=\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}},\quad i\geq 0.$ (16)
So, since $\xi_{j}<1$ and $q>1$, we have
$1>\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}}>\frac{1}{1+\frac{b(q-1)q^{i+2}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i+1}\xi_{j}}}\Rightarrow 1>a_{i}>a_{i+1},\quad i\geq 0.$ (17)
Furthermore, as $\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}<1$, we have
$0<x(t_{n})$ $=x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}a_{0}a_{1}a_{2}\dots a_{n-2}$
$<x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}(a_{0})^{n-1},$
leading to $\underset{n\to\infty}{\lim}x(t_{n})=0$.
Considering now (14), we can define $\tilde{a}_{i}$ as
$\tilde{a}_{i}=\frac{1}{\xi_{i+1}}\cdot\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}},\quad i\geq 0.$ (18)
From the previous case, we already know that
$\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}},\quad i\geq 0,$
tends to zero as $i$ grows. Now we prove that $\xi_{i+1}<\xi_{i}<1$ for all $i\geq 0$. For a certain $i\geq 0$, suppose that $\xi_{i+1}<\xi_{i}$. Then,
$\frac{1+c(q-1)q^{i+1}t_{0}}{1+b(q-1)q^{i+1}t_{0}}<\frac{1+c(q-1)q^{i}t_{0}}{1+b(q-1)q^{i}t_{0}},$
which is equivalent to
$b(q-1)q^{i}t_{0}-b(q-1)q^{i+1}t_{0}+c(q-1)q^{i+1}t_{0}-c(q-1)q^{i}t_{0}<0$
$\Leftrightarrow(c-b)(q-1)^{2}q^{i}t_{0}<0.$
This is always true since, by definition, $b>c$ and $q>1$. This means that $\xi_{i+1}<\xi_{i}$ for all $i\geq 0$. Combined with (15) and (17), this leads to
$\frac{1}{\xi_{i+1}}\cdot\frac{1}{1+\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}}<\frac{1}{\xi_{i+2}}\cdot\frac{1}{1+\frac{b(q-1)q^{i+2}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i+1}\xi_{j}}}<1\Rightarrow\tilde{a}_{i}<\tilde{a}_{i+1}<1,\quad i\geq 0.$
Therefore,
$0<y(t_{n})$ $=\frac{y(t_{0})}{\xi_{0}}\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\tilde{a}_{0}\tilde{a}_{1}\tilde{a}_{2}\dots\tilde{a}_{n-2}$
$<\frac{y(t_{0})}{\xi_{0}}\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}(\tilde{a}_{n-2})^{n-1},$
leading to $\underset{n\to\infty}{\lim}y(t_{n})=0$. Also, $\underset{n\to\infty}{\lim}z(t_{n})=N$. Thus, all solutions of (5) converge to the equilibrium point $(0,0,N)$ under the condition $b>c$. ∎
To address the case when $\mathcal{R}_{0}<1$, we make use of the following lemma.
Lemma 1.
Let $\mathcal{R}_{0}<1$, that is, $b<c$. Then, $\xi_{j}>1$ for all $j\geq 0$. Moreover, considering (16), then $a_{i}>a_{i+1}$ for all $i\geq 0$.
Proof.
Let us consider that $a_{i}>a_{i+1},\forall i\geq 0$. This means that
$\frac{b(q-1)q^{i+1}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}}<\frac{b(q-1)q^{i+2}t_{0}}{1+\bar{\kappa}\prod_{j=0}^{i+1}\xi_{j}},$
which is equivalent to
$q^{i+1}\left(1+\bar{\kappa}\displaystyle\prod_{j=0}^{i+1}\xi_{j}\right)<q^{i+2}\left(1+\bar{\kappa}\displaystyle\prod_{j=0}^{i}\xi_{j}\right)\Leftrightarrow(1-q)+\bar{\kappa}\prod_{j=0}^{i}\xi_{j}(\xi_{i+1}-q)<0.$
Since $q>1$, the previous relation is always true if $\xi_{i+1}-q<0$, that is,
$\frac{1+c(q-1)q^{i+1}t_{0}}{1+b(q-1)q^{i+1}t_{0}}<q.$
Simplifying, we obtain that
$1+(c-bq)(q-1)q^{i+1}t_{0}-q<0,$ (19)
which is always true provided
$c-bq<0\Leftrightarrow q>\frac{c}{b}.$
On the other hand, (19) is equivalent to
$(1-q)(1-(c-bq)q^{i+1}t_{0})<0.$ (20)
Since $q>1$, (20) holds true if and only if
$(1-(c-bq)q^{i+1}t_{0})>0,$
which is equivalent to
$\frac{1}{(c-bq)q^{i+1}}>t_{0}>0.$ (21)
Finally, (21) is true if and only if
$c-bq>0\Leftrightarrow q<\frac{c}{b}.$
If $q=\frac{c}{b}$, from (20) we get
$1-\frac{c}{b}<0.$
Since, by hypothesis, $\mathcal{R}_{0}<1$, then $\frac{c}{b}>1$, and thus, the result holds for any $q>1$. ∎
Theorem 3.
Let $\mathcal{R}_{0}<1$. Then, under the conditions of Lemma 1, all solutions (5) of system (4) converge to the equilibrium $(\alpha,0,N-\alpha)$ for some $\alpha\in\,]0,N]$.
Proof.
Since $\mathcal{R}_{0}<1$, that is, $b<c$, then $\xi_{j}>1$. Note that considering (16), it is clear that $a_{i}<1,\forall i\geq 0$. From Lemma 1, we have $1>a_{i}>a_{i+1}$, for any $q>1$. So,
$x(t_{n})$ $=x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}a_{i}>x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}a_{n-2}$
$\Rightarrow x(t_{n})$ $>x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}(a_{n-2})^{n-1}.$
Since $(a_{n-2})^{n-1}\longrightarrow 0$ as $n\to\infty$, then $x(t_{n})=x(t_{0})\frac{\bar{\kappa}+1}{\bar{\kappa}+1+b(q-1)t_{0}}\displaystyle\prod_{i=0}^{n-2}a_{i}=\alpha$, meaning that $\underset{n\to\infty}{\lim}x(t_{n})=\alpha$, with $\alpha\in\,]0,N]$. Having this, and recalling $\tilde{a}_{i}$ from (18), it just remains left to verify that $\prod_{i=0}^{n-2}\frac{1}{\xi_{i+1}}\longrightarrow 0$. This is clearly true since in this case $\xi_{i+1}>1$ for all $i\geq 0$. Therefore, $\underset{n\to\infty}{\lim}y(t_{n})=0$. It is now trivial to conclude that $\underset{n\to\infty}{\lim}z(t_{n})=N-\alpha$. From what we have shown, all solutions (5) converge to the equilibrium $(\alpha,0,N-\alpha)$. ∎
3 Examples
In this section, we present some illustrative examples confronting systems (1) and (3) under various parameter values. We start by plotting in Figure 1 the case where $b>c$, in which can be observed that the solution converges to the equilibrium point $(0,0,N)$ in both continuous and quantum cases.
(a) Solution of (1)
(b) Solution of (4)
Figure 1: Susceptible, Infected and Removed individuals with $b=0.3$, $c=0.1$, $x(t_{0})=0.6$, $y(t_{0})=0.4$, $z(t_{0})=0$, $q=1.1$, and $t_{0}=0.01$.
In contrast, Figure 2 examines the scenario where the parameters satisfy $b<c$. In this case, it is observable that the solution converges to the equilibrium ($\alpha,0,N-\alpha$), where $\alpha\in\mathbb{R}^{+}$. These examples align with the results obtained in Section 2, reinforcing consistency between continuous and quantum SIR models.
(a) Solution of (1)
(b) Solution of (4)
Figure 2: Susceptible, Infected and Removed individuals with $b=0.3$, $c=0.6$, $x(t_{0})=0.6$, $y(t_{0})=0.4$, $z(t_{0})=0$, $q=1.1$, and $t_{0}=0.01$.
As expected, for $q\approx 1$, the obtained quantum solutions closely approximate the solutions of the classical continuous SIR model (see Figures 1 and 2). Note that the parameter values and initial conditions used in the example shown in Figure 1 are consistent with condition (15).
For higher values of $q$, in the specific case where $R_{0}<1$, the susceptible population $x$ continues to converge to a certain $\alpha\in\,]0,N]$, although to an increasingly smaller positive value (see Figure 3). This last result is in line with Lemma 1 and Theorem 3.
Figure 3: Susceptible individuals for different values of $q$ and $b<c$.
4 Conclusion
In this work, we proposed a quantum SIR model considering the concept of $q$-derivative and a particular rule of Mickens’ method. The equilibrium points were found, and the non-negativity and boundedness of solutions were proved. We derived the exact solution of the quantum SIR model and proved asymptotic stability. We end providing some examples that illustrate the exact solution obtained.
While the assumption of constant rates is often not realistic in practice, having an exact solution for such a case can serve as a benchmark for comparing the behavior of the system under different parameter values. Our examples show that, for a $q\rightarrow 1^{+}$, our system behaves similarly to the continuous model. Moreover, our results highlight the potential of quantum time as an alternative framework for discrete modeling in epidemiology. Future work may include extending this approach to more complex compartmental models, as well as exploring its applications in other scientific domains.
Declarations
Conflict of interest
The authors declare no conflicts of interest.
Data availability
The manuscript has no associated data.
Funding
The authors were partially supported by the Portuguese Foundation for Science and Technology (FCT): Lemos-Silva and Torres through the Center for Research and Development in Mathematics and Applications (CIDMA) of University of Aveiro (UA), project UID/04106/2025 (https://doi.org/10.54499/UID/04106/2025); Vaz through the Center of Mathematics and Applications (CMA) of Universidade da Beira Interior (UBI), project UID/00212/2025 (https://doi.org/10.54499/UID/00212/2025). Lemos-Silva is also supported by the FCT PhD fellowship UI/BD/154853/2023 (https://doi.org/10.54499/UI/BD/154853/2023).
Acknowledgements
The authors are grateful to two reviewers for reading the initial submitted version carefully and for several comments and suggestions.